# Multi-modal Machine Learning Analysis of X-ray Absorption Near-Edge Spectra and Pair Distribution Functions: Performance and Interpretability towards Experimental Design

### v1

@@ 2.3 Average nearest-neighbor bond-length regression results
We saw similar results in the charge state prediction (see Fig. S10 and Table S5 in SI), where dPDF models outperformed the total-PDF models across all elements. Significant improvement over total-PDF was seen in Ti, Mn, and Cu, with F1 score increases in the 7-11% range, but this improvement was only modest for Fe (3%). For all metals, combining dPDF with XANES slightly improved the performance over using XANES alone. For most metals here (Ti, Mn, and Cu), as with bond length multimodal results, the dPDF features around the nearest-neighbor peak retained much more notable importance when combined with XANES. The only exception here is Fe, where replacing total-PDFs with dPDFs did not change the feature importance nor the multimodal model score significantly.
The single-modal and multimodal dPDF models were able to extract the charge state based on the correlation between charge state and nearest-neighbor bond lengths [brown1973empirical, brown1985bond]. Although the total-PDFs also contain this information, the models were more capable of extracting this information from the dPDFs for most metals here, further suggesting that dPDFs can be seen as a clearer signal for describing the coordination environment of a specific species.
@@ 3 Discussion & Conclusion
This work demonstrates how to use interpretable machine learning techniques in tandem with computationally generated multi-modal spectra to study the relative contributions of XANES and PDF to predictions of coordination number, charge (oxidation) state, and bond length. Feature importance plots reveal regions in the spectra from which the random forest models draw the most important insights and allow us to compare the relative utility of different spectra for the same prediction tasks in metal oxides. We anticipate that experimentalists who have access to comparable databases of reference spectra may be able to use this approach to compare the expressiveness of different modalities, both individually and in combination, without the need for a detailed forward model that can accommodate both spectral types. Assuming that simple ML models like random forests can capture the same trends that a more system-specific model would, this approach could also help plan experiments by assessing the value of incorporating an additional modality.
@@
For XANES, as previously observed we find that the pre-edge region is a strong predictor of coordination number for most metals investigated (Ti, Mn, and Cu), particularly for Ti where the pre-edge peak height and position are strongly dependent on Ti coordination number [farges_ti_1997]. This lends confidence to our approach. Pre-edge XANES also contains distinguishing information on Fe oxidation state [wilke2001oxidation]. Additionally, XANES edge energy [glatzel2009electronic, guda_understanding_2021] and peak positions [natoli1984distance] have been correlated with nearest-neighbor bond length, and this is reflected in the feature importance plots for most metals (Ti, Mn, and Cu), except for Fe. Although the post-edge region of XANES are less studied by traditional methods due to fitting difficulty [waychunas_synchrotron_1987-1], our ML models and others (such as [torrisi_random_2020, guda_understanding_2021]) found the post-edge XANES region to be highly predictive for most targets. On the other hand, PDF importance plots showed that the models were able to extract most information from nearest-neighbor PDF peaks. While coordination number and bond length can be extracted more directly from PDF peak heights and positions, the PDF importance plots confirmed that the charge state prediction is enabled by its correlation with the nearest-neighbor bond lengths [brown1973empirical, brown1985bond].
@@
XANES contains rich information about the absorbing atom’s electronic spectrum and has traditionally been used to obtain properties like charge state of the ions. It has been used much less to obtain detailed local structural information due to the nonlinearity of the absorption response, in contrast with absorption signal in the EXAFS region which is easily fit to the EXAFS equation[rehr2000theoretical]. However, ML techniques have increasingly enabled researchers to extract specific local structural information from XANES such as coordination environments and nearest-neighbor bond lengths for nanoparticles [timoshenko2017supervised, timoshenko2020silver] and transition metal oxides [carbone_classification_2019, torrisi_random_2020, zheng_random_2020], and aromaticity [tetef2021unsupervised] and chemical motifs [carbone2024accurate] of small organic molecules. And although generative AI was used to predict 3D structures of amorphous carbon [kwon2023spectroscopyguided] conditioned on XANES, overall, the scope of what can be extracted from XANES is still somewhat limited to specific prediction targets and simpler chemical systems, so we might expect structural probes such as PDF to outperform XANES for purely structural information.
@@
However, in our work, XANES models outperformed total-PDF models for all metals and prediction tasks, including geometrical properties like coordination number and bond length. This adds to the growing body of work [timoshenko2019inverting, penfold2024machinelearning] on extracting precise local structural information directly from XANES. The intrinsic correlation between electronic and atomic structure in XANES evidently allows random forest models to effectively leverage these relationships. Our findings also raise the question of why XANES outperformed PDF in most cases. It could be that the structural information is encoded more efficiently in XANES; for example, the non-linear relationship between the electronic spectrum and the structure could increase sensitivity.
@@
This underscores the potential utility of developing methods that can extract differential PDFs experimentally, such as using isotopic substitution of neutrons [soper1977neutron, enderby1995ion, zhao1998neutron] or anomalous differential diffraction methods [aur1983local, kofalt1986differential, waseda1984novel, waseda2002anomalous, petkovApplicationAtomicPair2000, petkovApplicationDifferentialResonant2018a]. These methods exist but are currently difficult and not widely used. Another useful extension of this work would be to do the same analysis comparing XANES and EXAFS input signals, or PDF and EXAFS.
As a caveat we note that, like any study drawing conclusions from ML, our results are tied to the dataset that was used. We are more confident that the observed trends would apply for similar systems (bulk ordered transition metal oxide crystals), and caution that extrapolation to other classes of system could and should be verified with comparable libraries of structures and spectra (surface structures, nanoparticles, disordered materials, non-oxide systems, and so on). Additionally, intermediate featurization techniques could potentially improve the model performance described in this study, but we leave this exploration for future work.
@@ 4.1 Datasets
Our datasets consist of transition metal oxide structures available on the Materials Project (MP) database [jain2013commentary]. We curated four separate datasets, one each for the four metal elements: Ti, Mn, Fe, and Cu. These metals had the most MP structures available that met the requirements of our prediction tasks. Depending on the targets (charge state, coordination number, and bond length), the sample size varied between $\approx 600$ and 1300 compounds (see Tables S1, S2, and S3 in SI). Each dataset included the Pymatgen-compatible structures, calculated K-edge XANES spectra with the transition metal as the absorbing atom, and total- and differential-PDFs computed by us.
The XANES spectra available on MP were computed via FEFF calculations [mathew2018high]. We used Diffpy-CMI [billinge2023atomic, juhas;aca15] to compute the total- and differential-PDFs of all structures in our datasets. In detail, we used Diffpy-CMI ’s PDFCalculator module to compute both the total- and differential-PDFs of the structures from their crystallographic coordinates. We first calculated the total-PDFs in the range $0\leq r\leq 10$ and interpolated onto a 100-pt Nyquist-Shannon grid [farrow2011nyquist] assuming measured PDFs with $q_{max}=30$-1. We did the same for the differential-PDFs, with the additional step of filtering for the species of interest (the transition metal in our case) when using PDFCalculator. We used unscaled PDFs and XANES because they gave slightly better model performances than when using scaled inputs. Previously, Torrisi et al. found that scaling did not significantly change pointwise model interpretability trends [torrisi_random_2020].
@@
Next, we extracted target quantities (charge state, coordination number, and mean nearest-neighbor bond length) from the structures as the ground truths for supervised learning. Charge state was extracted using Pymatgen’s BVAnalyzer module. To avoid ambiguities from averaging variable charge states across different sites, we further subsampled our dataset for this task and kept only structures in which all transition metal cations had the same oxidation state. This makes it easier to understand the ability of random forest models to extract charge state trends as this simplifies the problem space. Further, to make sure there is sufficient representation for each class and equal number of classes in all datasets (three classes), we included only Ti, Mn, and Fe oxides with metal cation oxidation states of 2+, 3+, and 4+, and Cu oxides with charge states 1+, 2+, and 3+.
Metal coordination numbers were obtained using Pymatgen’s CrystalNN module, which uses Voronoi decomposition and solid angle weights to calculate the most probable coordination environment around a specified atom [pan2021benchmark]. We first extracted the coordination number (number of nearest neighbors) of each metal site in the structure. Similar to our charge state dataset curation, we filtered for structures in which all of the metal sites had the same coordination number and only included those with coordination numbers of 4, 5, or 6 in the prediction task. This resulted in datasets of size $N\approx$900 for each transition metal element used here.
@@ 4.2 Model training
We trained random forest classifiers for charge state and coordination number prediction and random forest regressors for mean nearest-neighbor bond length prediction. Both model types made use of scikit-learn [pedregosa2011scikit]. For all targets, we held out 20% of samples from each dataset for testing and trained on the remaining 80%. Initially, three models were trained for each transition metal element, with varying inputs: one model with total-PDF as input, another with XANES as input, and the third with both XANES and total-PDF as input. We used F1 scores (class-wise and weighted mean) as the performance metric for the classification tasks and Root Mean Square Error (RMSE) for the regression task. We report the RMSE as the percentage of mean bond length to contextualize the magnitude of the model error relative to the true values.
@@
For charge state and coordination number, we also computed baseline F1 scores for comparison with our model scores. Our datasets for these classification tasks consist of three imbalanced classes, often with a clear majority class, so we computed the baseline F1 scores based on a trivial classifier that labels all samples as the majority class. The baseline score depends on the class proportions, so it is unique for each metal and each target.
We also estimated baseline RMSEs for the mean nearest-neighbor bond length prediction. Our “naive” model in this case estimated the mean bond length by extracting the the location (in Å) of the nearest-neighbor peak from the total-PDF using the peak-finding algorithm in scipy [virtanen2020scipy]. The RMSEs of these predictions were much higher than the random forest model test errors so they were not shown in Fig. 3. The baseline errors ranged from 14-20.7% of the mean bond length, much larger than 2-4% from the random forest models.
5 Data availability
@@
Acknowledgements and funding information
We thank Weike Ye and Linda Hung for helpful discussions. This work was funded by Toyota Research Institute, grant number PO-002332.

### v2

@@ 2.3 Mean nearest-neighbor bond-length regression
We saw similar results in the oxidation state prediction (see Fig. S10 and Table S6 in SI), where dPDF models outperformed the total-PDF models across all elements. Significant improvement over total-PDF was seen in Ti, Mn, and Cu, with F1 score increases in the 7-11% range, but this improvement was only modest for Fe (3%). For all metals, combining dPDF with XANES slightly improved the performance over using XANES alone. For most metals here (Ti, Mn, and Cu), as with bond length multimodal results, the dPDF features around the nearest-neighbor peak retained much more notable importance when combined with XANES. The only exception here is Fe, where replacing total-PDFs with dPDFs did not change the feature importance nor the multimodal model score significantly.
The single-modal and multimodal dPDF models were able to extract the oxidation state based on the correlation between oxidation state and nearest-neighbor bond lengths [46, 47]. Although the total-PDFs also contain this information, the models were more capable of extracting this information from the dPDFs for most metals here, further suggesting that dPDFs can be seen as a clearer signal for describing the coordination environment of a specific species.
@@ 3 Discussion
This work demonstrates how to use interpretable machine learning techniques in tandem with computationally generated multimodal spectra to study the relative contributions of XANES and PDF to predictions of coordination number, oxidation state, and bond length. Feature importance plots reveal regions in the spectra from which the random forest models draw the most important insights and allow us to compare the relative utility of different spectra for the same prediction tasks in metal oxides. We anticipate that experimentalists who have access to comparable databases of reference spectra may be able to use this approach to compare the expressiveness of different modalities, both individually and in combination, without the need for a detailed forward model that can accommodate both spectral types. Assuming that simple ML models like random forests can capture the same trends that a more system-specific model would, this approach could also help plan experiments by assessing the value of incorporating an additional modality.
@@
For XANES, as previously observed we find that the pre-edge region is a strong predictor of coordination number for most metals investigated (Ti, Mn, and Cu), particularly for Ti where the pre-edge peak height and position are strongly dependent on Ti coordination number [52]. This lends confidence to our approach. Pre-edge XANES also contains distinguishing information on Fe oxidation state [45]. Additionally, XANES edge energy [60, 28] and peak positions [61] have been correlated with nearest-neighbor bond length, and this is reflected in the feature importance plots for most metals (Ti, Mn, and Cu), except for Fe. Although the post-edge region of XANES are less studied by traditional methods due to fitting difficulty [50], our ML models and others (such as [24, 28]) found the post-edge XANES region to be highly predictive for most targets. On the other hand, PDF importance plots showed that the models were able to extract most information from nearest-neighbor PDF peaks. While coordination number and bond length can be extracted more directly from PDF peak heights and positions, the PDF importance plots confirmed that the oxidation state prediction is enabled by its correlation with the nearest-neighbor bond lengths [46, 47].
To compare the relative importance of the pre-edge and post-edge regions of XANES, we also tried training the models on each region separately (see results in Sec. 6 in SI) for the coordination number prediction. For all four metals, including Ti where the pre-edge importance is the most prominent, the models trained on post-edge XANES performed better than those trained on the pre-edge region alone, but still slightly worse than when using the whole spectrum as input (see Fig. S13 in SI). The results are consistent with the findings in [22], where their CNN models for coordination number prediction performed worse when trained on the pre-edge only compared to the whole spectrum, indicating that the post–edge region contains important information. Additionally, we combined each region of XANES with PDFs and dPDFs and re-trained the models. While the test scores did not change significantly whether we used the pre-edge, post-edge, or the whole range of XANES in the combined models (see Figs. S14 and S16 in SI), the feature importance plots (Figs. S15 and S17 in SI) highlight that Ti is the only metal here where both pre-edge and post-edge XANES heavily dominated the prediction in the presence of PDFs/dPDFs. For other metals (Mn, Fe, and Cu), post-edge XANES and PDF/dPDF features are closer in importance and contain more coordination number information than pre-edge XANES.
@@
XANES contains rich information about the absorbing atom’s electronic spectrum and has traditionally been used to obtain properties like oxidation state of the ions. It has been used much less to obtain detailed local structural information due to the nonlinearity of the absorption response, in contrast with the absorption signal in the EXAFS region which is easily fit to the EXAFS equation [62]. However, ML techniques have increasingly enabled researchers to extract specific local structural information from XANES such as coordination environments and nearest-neighbor bond lengths for nanoparticles [21, 63] and transition metal oxides [22, 24, 23], and aromaticity [26] and chemical motifs [42] of small organic molecules. And although generative AI was used to predict 3D structures of amorphous carbon [25] conditioned on XANES, overall, the scope of what can be extracted from XANES is still somewhat limited to specific prediction targets and simpler chemical systems, so we might expect structural probes such as PDF to outperform XANES for purely structural information.
@@
However, in our work, XANES models outperformed total-PDF models for all metals and prediction tasks, including geometrical properties like coordination number and bond length. This adds to the growing body of work [17, 18] on extracting precise local structural information directly from XANES. The intrinsic correlation between electronic and atomic structure in XANES evidently allows random forest models to effectively leverage these relationships. Our findings also raise the question of why XANES outperformed PDF in most cases. It could be that the structural information is encoded more efficiently in XANES; for example, the non-linear relationship between the electronic spectrum and the structure could increase sensitivity. In our work, the ML models are able to learn a direct mapping from the structural signal to the electronic properties, and from the electronic signal to the structural properties.
@@
This underscores the potential utility of developing methods that can extract differential PDFs experimentally, such as using isotopic substitution of neutrons [64, 65, 66] or anomalous differential diffraction methods [54, 55, 56, 57, 67, 59]. These methods exist but are currently difficult and not widely used. One reason is that the experiments and the subsequent data analysis are difficult. A second limitation of the anomalous differential PDF is that the $Q$-range available from these measurements, and therefore the real-space resolution in the PDF, is limited by the energy of the incident x-rays. For example, first-row transition metals such as studied here the absorption edges are below 10 keV and the accessible $Q$-range in an adPDF measurement would be well below $Q_{\mathrm{max}}=10$-1. An interesting future study would explore in more quantitative detail how low real-space resolution but chemically resolved adPDFs complement XANES.
Another useful extension of this work would be to do the same analysis comparing XANES and EXAFS input signals, or PDF and EXAFS.
As a caveat we note that, like any study drawing conclusions from ML, our results are tied to the dataset that was used. We are more confident that the observed trends would apply for similar systems (bulk ordered transition metal oxide crystals), and caution that extrapolation to other classes of system could and should be verified with comparable libraries of structures and spectra (surface structures, nanoparticles, disordered materials, non-oxide systems, and so on). It is also important to note discrepancies between the theoretical spectra used in our study and the experimental data. For example, experimental PDFs might have lower real-space resolution due to smaller $Q_{\mathrm{max}}$, which depends on the X-ray energy and scattering angle (we assumed $Q_{\mathrm{max}}$ $=30$-1 in the calculation). For the case of XANES, FEFF calculations [14] are known to perform better in the post-edge region, where multiple scattering effects dominate, than in the pre-edge region, where electronic structure dominates. Another limitation is that we work with small defect-free ordered unit cells from the Materials Project, which represent a strongly idealized version of real materials. In real materials, the structures that are observed may differ for a variety of reasons, including but not limited to: spontaneous symmetry breaking, distortion into larger unit cells due to the Jahn-Teller effect, or displacive disorder due to finite temperature. Future applications of this method on individual systems of interest should attempt to include these effects as necessary in the reference structures that are used to generate input datasets. Additionally, intermediate featurization techniques could potentially improve the model performance described in this study, but we leave this exploration for future work.
@@ 4.1 Datasets
Our datasets consist of transition metal oxide structures available on the Materials Project (MP) database [16]. We curated four separate datasets, one each for the four metal elements: Ti, Mn, Fe, and Cu. These metals had the most MP structures available that met the requirements of our prediction tasks. In most of the oxide structures used here, the transition metal atoms’ nearest neighbors are all oxygen, but other species besides the metal and oxygen are present in some structures (see Table S4 in SI). Depending on the targets (oxidation state, coordination number, and bond length), the sample size varied between $\approx 600$ and 1300 compounds (see Tables S1, S2, and S3 in SI). Each dataset included the Pymatgen-compatible structures, calculated K-edge XANES spectra with the transition metal as the absorbing atom, and total- and differential-PDFs computed by us.
The XANES spectra available on MP were computed via FEFF calculations [15]. We used Diffpy-CMI [9, 6] to compute the total- and differential-PDFs of all structures in our datasets. In detail, we used Diffpy-CMI ’s PDFCalculator module to compute both the total- and differential-PDFs of the structures from their crystallographic coordinates. We first calculated the total-PDFs in the range $0\leq r\leq 10$ and interpolated onto a 100-pt Nyquist-Shannon grid [68] assuming measured PDFs with $Q_{max}=30$-1. We did the same for the differential-PDFs, with the additional step of filtering for the species of interest (the transition metal in our case) when using PDFCalculator. We used unscaled PDFs and XANES because they gave slightly better model performances than when using scaled inputs. Previously, Torrisi et al. found that scaling did not significantly change pointwise model interpretability trends [24].
@@
Next, we extracted target quantities (oxidation state, coordination number, and mean nearest-neighbor bond length) from the structures as the ground truths for supervised learning. Oxidation state was extracted using Pymatgen’s BVAnalyzer module. To avoid ambiguities from averaging variable oxidation states across different sites, we further subsampled our dataset for this task and kept only structures in which all transition metal cations had the same oxidation state. This makes it easier to understand the ability of random forest models to extract oxidation state trends as this simplifies the problem space. Further, to make sure there is sufficient representation for each class and equal number of classes in all datasets (three classes), we included only Ti, Mn, and Fe oxides with metal cation oxidation states of 2+, 3+, and 4+, and Cu oxides with oxidation states 1+, 2+, and 3+.
Metal coordination numbers were obtained using Pymatgen’s CrystalNN module, which uses Voronoi decomposition and solid angle weights to calculate the most probable coordination environment around a specified atom [69]. We first extracted the coordination number (number of nearest neighbors) of each metal site in the structure. Similar to our oxidation state dataset curation, we filtered for structures in which all of the metal sites had the same coordination number and only included those with coordination numbers of 4, 5, or 6 in the prediction task. This resulted in datasets of size $N\approx$900 for each transition metal element used here.
@@ 4.2 Model training
We trained random forest classifiers for oxidation state and coordination number prediction and random forest regressors for mean nearest-neighbor bond length prediction. Both model types made use of scikit-learn [70]. For all targets, we held out 20% of samples from each dataset for testing and trained on the remaining 80%. Initially, three models were trained for each transition metal element, with varying inputs: one model with total-PDF as input, another with XANES as input, and the third with both XANES and total-PDF as input. We used F1 scores (class-wise and weighted mean) as the performance metric for the classification tasks and Root Mean Square Error (RMSE) for the regression task. We report the RMSE as the percentage of mean bond length to contextualize the magnitude of the model error relative to the true values.
@@
For oxidation state and coordination number, we also computed baseline F1 scores for comparison with our model scores. Our datasets for these classification tasks consist of three imbalanced classes, often with a clear majority class, so we computed the baseline F1 scores based on a trivial classifier that labels all samples as the modal class. The baseline score depends on the class proportions, so it is unique for each metal and each target.
We also estimated baseline RMSEs for the mean nearest-neighbor bond length prediction. Our “naive” model in this case estimated the mean bond length by extracting the the location (in Å) of the nearest-neighbor peak from the total-PDF using the peak-finding algorithm in scipy [51]. The RMSEs of these predictions were much higher than the random forest model test errors so they were not shown in Fig. 3. The baseline errors ranged from 14-20.7% of the mean bond length, much larger than 2-4% from the random forest models.
Data availability
@@
Acknowledgements We thank Weike Ye and Linda Hung for helpful discussions. This work was funded by Toyota Research Institute, grant number PO-002332.